An object has a kinetic energy of 275 J and a momentum of magnitude 25.0 kg * m/s. Find the (a) speed and (b) mass of the object.
1
Expert's answer
2020-10-26T14:47:49-0400

p=25.0 kg * m/s

KE=275 J

"Ke=\frac{1}{2}mv^2"

"Ke=\frac{1}{2}mv^2⋅\frac{m}{m}"

"Ke=\frac{m^2v^2}{2m}=\frac{(mv)^2}{2m}"

"Ke=\frac{(mv)^2}{2m}"

mv=p

"Ke=\frac{(p)^2}{2m}"

"2m\u22c5Ke=p^2"

"m= \frac{p^2}{2Ke}=\frac{25^2}{2⋅275}=1.14 kg"


"p=mv"

"v=\frac{p}{m}=\frac{25}{1.14}=22\frac{m}{s}"